# Information Seeking Using AI Assistants

### v1

@@ Abstract
A good portion of a software practitioners’ day involves seeking and using information to support task completion. Although the information needs of software practitioners have been studied extensively, the impact of AI-assisted tools on their needs and information-seeking behaviors remains largely unexplored. To addresses this gap, we conducted a mixed-method study to understand AI-assisted information seeking behavior of practitioners and its impact on their perceived productivity and skill development. We found that developers are increasingly using AI tools to support their information seeking, citing increased efficiency as a key benefit. Our findings also amplify caveats that come with effectively using AI tools for information seeking, especially for learning and skill development, such as the importance of foundational developer knowledge that can guide and inform the information provided by AI tools. Our efforts have implications for the effective integration of AI tools into developer workflows as information retrieval and learning aids.
Index Terms:
AI-assisted tools, information seeking, software development, productivity
I Introduction
@@
Artificial Intelligence (AI) has revolutionized how we think about software engineering. According to the most recent StackOverflow Developer Survey, 76.7% of developers are already using or plan to incorporate AI-assisted software development assistants into their workflows [1]. Developers use AI-assisted software tools in a variety of contexts and to complete a diversity of tasks, most often hoping for increases in their productivity [2].
@@
A significant contributor to, or deterrent from, developer productivity is the time they spend seeking information to support the completion of their tasks [3, 4]. Over the years, we have acquired an in-depth understanding of the information needs developers have [5, 6, 7], the ways in which they attempt to meet their information needs [8], and the challenges they encounter in the process [9]. We also know that developers’ ability to meet information needs contributes to their ability to acquire and build expertise [10, 11].
@@
Given the changing landscape in developer tooling, many studies have investigated the use and impact of AI-assisted tools on developer engagement and productivity [12, 13, 14, 15] . However, most of these studies have focused on AI-assisted tool use in the context of understanding, generating, or modifying source code, providing little insight into the role AI plays in developer information seeking and its impact. To address this gap, we conducted a mixed methods study to answer the following research questions:
RQ1
@@
To answer our research questions, we first administered a survey to better understand to what extent developers are engaging with AI tools in this context. Building on those insights, we conducted a series of 17 interviews to further contextualize those experiences. Our findings reveal that while AI tools can offer increased efficiency when seeking information, there are caveats to realizing the benefits, such as avoiding over-reliance and building the necessary expertise to appropriately and effectively use AI tooling to meet information needs.
II Related Work
@@
The goal of our research is better understand information seeking with the availability of AI-assisted tools. Most relevant to our efforts investigations into the information needs and seeking behaviors of developers and human-centric concerns regarding the use of AI tools in software development.
II-A Information Needs of Developers
@@
Previous studies have examined the information needs of developers during the software development process. One of the first efforts centered on developer information needs was conducted by Ko et al. [5], where they observed developers in their work to better understand information needs of collocated software teams. Also aiming to gain a broad understanding of developer information needs is the work done by Fritz et al. [8], who conducted a series of 11 interviews to explore the use of information fragments to answer developers’ questions. On the flip side, LaToza et al. [9] surveyed 179 professional developers regarding hard-to-answer questions about code to better understand information needs developers encounter challenges trying to fulfill.
@@
Some prior efforts are more focused, with an interest in understanding more specific information needs. Breu et al. [16] analyzed information needs in bug reports by examining questions asked in six hundred bug reports from the Mozilla and Eclipse projects. Liu et al. [6] investigated API-related developer information needs on StackOverflow by annotating StackOverflow questions to APIs. Phillips et al. [7] explored information needs for integration decisions in the release process of large-scale parallel development by interviewing seven release managers. With the emergence of AI-assisted tools, our builds on prior efforts by providing insights into the ways in which these tools have changed software practitioners’ information-seeking behaviors.
II-B AI Tool Use in Software Engineering
Researchers have also studied the impact of integrating AI tools into software engineering processes. Zheng et al. [17] conducted a literature review and identified 123 papers focusing on large language models (LLMs) in software engineering, highlighting their utility for various development tasks. Most other works in human-centered AI-assisted tool use focus on understanding and supporting developer interactions. Barke et al. [18] investigated how developers interact with AI programming assistants like GitHub Copilot. They classified interaction modes into two types: acceleration mode, where developers know the next steps and use Copilot to speed up their work, and exploration mode, where developers are unsure and use Copilot to discover possible solutions. Liang et al. [19] conducted a large-scale survey to investigate the usability of AI programming assistants, focusing on the successes and challenges developers face when using these tools. Johnson et al. [20] developed the PICSE framework, which identifies key factors influencing engineers’ trust and usage of traditional and AI-assisted software tools. Their framework was derived from interviews with software practitioners. Given the concern regarding the impact of AI-assisted tool use on productivity, prior work has also investigated how we can measure developer productivity when using code completion tools like GitHub Copilot1[12]. Our work builds on these prior efforts by providing insights into how these developer interactions with AI tools influence not just task completion, but also their information-seeking behaviors and its impact on their productivity and skill development.
III Methodology
@@
III-A The Survey
We designed a 20-minute survey administered through Qualtrics to engage developers in our research.  [^1] This research is approved under IRBNet #: [Blind for review]. Our survey was divided into four sections. The first section asked a series of demographic and background questions, such as current job role and years of programming experience. The design of this section was heavily influenced by the annual StackOverflow Developer Survey [21]. The next section asked questions about the specific AI-assisted tools they use, the frequency and purpose of use, and rationale behind their usage. We also asked questions about their information seeking behaviors with and without AI assistance, including the advantages and disadvantages for each. To assess the impact of AI tools on development tasks, we also asked respondents to evaluate how these tools influenced their approach to tasks like debugging, testing, implementing, and planning. In the final section of the survey, we included questions regarding the role of AI tool in learning and integrating new technologies, such as the impact on learning curves and the potential pitfalls. We also gave respondents the opportunity at the end of the survey to express interest in a follow-up interview (Section III-B). Prior to administering our survey, we piloted with 3 respondents. This helped ensure clarity and identify any issues before deployment. We excluded this data from our final dataset. The survey instrument is available in our supplemental materials.
@@ III-A The Survey
Given we distributed our survey through social platforms, there is a heightened risk of receiving invalid responses [22]. To mitigate the potential for analyzing invalid data, we filtered out responses that were completed in under three minutes or incomplete. This initial filtering left us with 168 responses. We further excluded responses that contained irrelevant content in their open-ended answers, as this suggests they may not have been giving due consideration to the survey. This resulted in our final dataset, containing 128 valid responses, which we used to report our findings.
TABLE I: Age Groups of Survey Respondents
@@
TABLE II: Professional Programming Experience
@@
Active UsersOccasional Users002020202040404040606060608080808010010010010083.683.683.683.616.416.416.416.4Percentage (%)
Figure 1: Usage of AI Tools Among Survey Respondents
TABLE III: Job Titles of Survey Respondents
@@
TABLE IV: AI-Assisted Tools Used by Survey Respondents
@@
In our final dataset of responses (128), the age of our survey respondents ranged from 18 to over 65, median being 25-34 [I] with of 4 years of experience [II]. Prior research suggests that developers with less experience spend more time seeking information and learning as they transition from novices to experts [10] [11] [23], making this demographic particularly valuable for our study. Majority of our survey respondents actively use AI tools [1] such as ChatGPT and GitHub Copilot [IV] in their work. In addition to formal education, a majority of our survey respondents also used books and online resources such as forums and courses to learn programming. We interviewed a subset of these respondents, depending on their interest and availability and report on their demographics in  V.
@@
To analyze our survey, we first mapped each survey question to our research questions. We categorized experience levels of participants into five different groups, as outlined in Table II. For closed ended questions, we ran descriptive statistics and applied Fisher’s exact test with Bonferroni correction to identify any correlations. We also report on frequencies to supplement our findings. For open-ended responses, we used thematic summaries to report findings due to the small sample size of valid and useful responses. In the following section, we only report significant findings.
III-B The Interviews
To supplement our survey findings, we conducted interviews over Zoom to gather detailed insights into participants’ information-seeking practices, experiences with AI-assisted tools, and their impact on productivity and skill development. This helped us obtain a variety of perspectives and allowed for immediate follow-up questions for clarification. Each interview began with a discussion of participants’ background in software development, their experiences with AI-assisted tools, and how these tools fit into their workflow. Topics covered included the impact of AI tools on productivity, skill development, and problem-solving strategies. We also asked questions regarding their roles and responsibilities, the typical activities in their jobs, and their reliance on other developers or software tools. This includes what they do use, or avoid use of, AI tools for when information seeking and their satisfaction with the interactions. We elicited specific examples of their use to better understand aspects such as workflow integration and problem solving approaches. Lastly, we inquired about the impact team dynamics and productivity.
We piloted the interview protocol following best practices with one author and two participants to identify and correct any issues. The final interview protocol is available in our supplemental materials. We continued interviewing participants until we reached theoretical saturation, where no new themes or insights emerged from the interview [24]. Table V details the background of our interview participants.
TABLE V: Interview Participant Background
@@ III-B The Interviews
To analyze our interview data, we used qualitative coding and thematic analysis, beginning with the creation of an initial codebook based on questions asked during the participant interviews. In the first iteration, all authors independently labeled an interview transcript using the initial codebook and then collectively discussed their findings, which led to the emergence of a new set of codes. After agreement on the codes and how they should be used, we incorporated the emergent codes to finalize the codebook. Following this, the first author applied the final codebook to label the remaining transcripts.
To ensure the rigor and validity of our efforts, we invited an external auditor to review a subset of raw coded segments from the interview transcripts along with the codebook. We asked the auditor to confirm that the inferences we made from the data were sound. We discussed and clarified any concerns and updated code descriptions when deemed necessary. After validating the codes, the first author reviewed all coded segments under each code and documented the overarching themes as they emerged. Finally, we mapped the codes to our research questions and iteratively identified any overlapping themes. Given our goal was to uncover themes present in the data, rather than their frequency,and following best practices for interview analysis [25], we do not report any Inter-rater Reliability (IRR).
IV Results
The goal of our efforts are to uncover when, why, and how developers are using AI tool to support their information seeking. Below we discuss our findings regarding the considerations developers may make when deciding to engage with AI tools to meet their information needs.
IV-A Information Seeking with AI tools (RQ1)
@@
Our findings suggest that developers are often using AI tools to support their information seeking (Figure 2 and Figure 3), with the majority reporting using AI tools at least half of the time in their work for support (104) and seeking information (93). However, according to the experiences reported in our interviews, their goals when doing so are most often broad and aimed at understanding the necessary considerations for completing their task.
5 (Always)4321 (Never)002020202040404040414141413131313132323232171717171111# Participants
Figure 2: Frequency of AI Use for Support
@@
Figure 3: Frequency of AI Use for Information Seeking
According to our survey respondents, the most notable benefits of using AI tools for information seeking include increased efficiency (92), more personalized responses (75), and programming language-agnostic support (43). One of our interview participants noted that using AI tools requires active engagement since “the AI is driving it” and this can go “on a tangent or in the right direction”. A few participants also discussed how using AI tools helps reduce the “burden” imposed on other developers when asking for help, as they then have to “spend time on our issues”.
Most participants reported relying on AI tools when trying to understand best practices, discover new libraries or solutions or explore trade-offs between different libraries and implementations. In this context, one participant stated “even if it can’t solve a complex coding problem, it can often point me to best practices for how to solve that problem.”[P14] Many participants noted that AI tools can help find links to relevant tutorials for in-depth learning while some participants reported that it can provide tailored explanations for new concepts using analogies or equivalents from their existing knowledge.
“And then I would also ask the GPTs or AI assistant for relevant links or relevant tutorials, so that I can actually learn about this process in general, and then apply that methodology or principle in my specific work.”[P8]
Participants in our study also cited AI tools as being useful for providing clear answers to generic conceptual questions, which can also aid in identifying keywords for further online searching and validation. Another common use case among participants when information seeking is to use AI tools to recall previous knowledge or explain code. Participants mentioned value in using AI tools for synthesizing relevant information from documentation, which includes providing boilerplate code, pinpointing highly specific issues or information. One participant shared a concrete example of this
@@
UX Issues
While participants in our study outlined numerous benefits to AI-assisted information seeking, concerns surfaced regarding the usability of AI tools. Most prominent were discussions regarding the ways in which the AI tool communicate the information. A major theme that emerged was around the conversational tone used by most AI-assisted tools. One participant discussed how this can undermine the utility of AI tools for tasks like code reviews.
“But for something like code review, it’s a bit riskier because error handling is not just a nice-to-have; it’s a must-have. While the AI correctly identifies many issues, it frames them as improvements rather than necessities. For example, it might suggest adding error handling or not hardcoding secrets, or it might recommend CSRF protection on endpoints as good improvements. However, these practices are actually mandatory.”[P15]
In addition, some participants cited that the tendency for AI tools to agree with the user also decreases their reliability as an information seeking tool. When recalling their own experiences, one participant stated, ”It provided the correct information initially but later changed its stance to align with what I was saying. So now I don’t know when to trust it.” Participants also reported issues with inappropriate information framing, finding the level of detail present in their responses to be ineffective, noting that “Sometimes they tend to give out too much information, sometimes too little.” We discuss more about the impact this has on information seeking when trying to learn new technologies in Section IV-C.
@@
Figure 4: Traditional vs. AI Preference
11footnotetext: Legend for task symbols: D = Debugging, CA = Code Analysis, R = Refactoring, $C$ = Code Comprehension, B = Brainstorming, LNT = Learning New Technologies.
Validation Techniques
While our findings point to consistent use of AI tools for information seeking, we found participants often only partially rely on information provided by AI tools, acknowledging the limitations of AI-generated solutions. Expressing sentiment shared by many participants, one participant stated, “Chat GPT 3.5 was providing a poor and incorrect implementation for network config files that were in YAML. ChatGPT3.5 also provided a partial answer that it scraped from Python docs regarding ArgParse library. I had to go to the actual docs to find it had a lot of missing context and prior steps in order to use the code it generated.” Prior work has emphasized concerns regarding the accuracy, privacy and security issues that can come with using AI tools  [20]. In addition to these concerns, our survey respondents indicated that when using AI tools to seek information they have concerns regarding decreases in self-directed learning and skill-building(76), privacy and security (66), and reduced information accuracy (60).
Given the skepticism around AI-generated information, our participants also discussed a variety of methods for validating the information provided. When it comes to information seeking endeavors that result in generated code or information used to modify their code, one common approach among our participants was unit testing. This involves running unit tests against any code modifications resulting from AI tool use to verify its functionality, check the output against expected results, and identify any issues or bugs.
Some participants noted using AI tools to support their validation efforts. In some cases, this meant using cross-validation, where they would compare responses from multiple AI assistants to identify conflicts or inconsistencies. For example, one participant stated that when one AI tool generates some code that needs an exhaustive set of test cases, “it’s especially helpful if you use multiple AI assistants for this” as it will help them “find a whole pile of issues”.
A common approach regardless of code being provided was iterative refinement, an AI-supported strategy where participants provide feedback to the AI assistant when the generated information is not satisfactory, modify prompts to obtain more accurate or appropriate information, and continuously refine the process until the desired outcome is achieved. They also rely on their own research and expertise, along with the expertise of others, to validate the information and confirm the accuracy and reliability of the AI-generated solutions. Or, as put by one participant, “there is always Google”. Another participant shared an experience they had with this kind of iterative solution refinement:
“I was trying to solve an error. And no matter how much I changed the point or how much I explained, it was still giving me the same solution. So I had to change my approach. I just went over to Google, and I entered whatever information I needed for that particular library to run. I went to StackOverflow, asked for different solutions and one of those solutions worked for me.”[P12]
@@
The most common resource mentioned for information seeking without AI assistance was StackOverflow, which participants cited using for context-specific tasks, such as determining the appropriate regular expressions for their use case or environment setup, in the hopes that other developers have encountered similar scenarios. Another common source participants mentioned using was Reddit, where they would seek general guidance, such as tool recommendations, from specific subreddits. We also found that when it comes to tasks that require specific information, knowledge, or context, or when working with specialized or niche libraries, our participants sought support directly from other developers or support teams on contract. When attempting to summarize the capabilities of AI tools for supporting information seeking in comparison to working with other developers, one participant stated, “It does it about as well as I would expect for a junior software engineer with a little bit more domain knowledge than I would expect from a a junior software engineer, but without a lot of the personality or experience, or trust that I would get from a junior software engineer.”
@@
Regardless of the benefits, and perhaps due in part to the concerns, that come with using AI tools for seeking information, participants in our study highlighted several external factors that influence the impact and their use of AI tools in software development. One such factor is organizational restrictions that aim to reduce the risks associated with using AI tools, which may include limitations on using certain AI tools like Copilot or the public version of ChatGPT. Our findings suggest organizations may also have agreements with providers such as Microsoft , driven by concerns regarding protection of Personally Identifiable Information (PII). Legal protections offered by major AI providers like Microsoft and Google, which pledge to defend users in AI-related copyright lawsuits, may also influence the confidence and extent to which organizations adopt AI tools.
Participants also noted that their usage of AI tools for information seeking is heavily influenced by the adoption and promotion of AI within the organization. Some organizations are proactive in adopting and promoting AI use, aiming to utilize these tools responsibly and in a controlled manner, while others feel “it’s not worth your time”. However, this proactive stance can sometimes lead to an overshadowing of traditional engineering practices, resulting in less investment in non-AI related professional development.

### v2

# The Evolution of Information Seeking in Software Development: Understanding the Role and Impact of AI Assistants

@@ Abstract.
About 32% of a software practitioners’ day involves seeking and using information to support task completion. Although the information needs of software practitioners have been studied extensively, the impact of AI-assisted tools on their needs and information-seeking behaviors remains largely unexplored. To addresses this gap, we conducted a mixed-method study to understand AI-assisted information seeking behavior of practitioners and its impact on their perceived productivity and skill development. We found that developers are increasingly using AI tools to support their information seeking, citing increased efficiency as a key benefit. Our findings also amplify caveats that come with effectively using AI tools for information seeking, especially for learning and skill development, such as the importance of foundational developer knowledge that can guide and inform the information provided by AI tools. Our efforts have implications for the effective integration of AI tools into developer workflows as information retrieval systems and learning aids.
software engineering, developers, mixed-methods study, ai, productivity, information seeking, skill building
journalyear: 2025copyright: ccconference: 33rd ACM International Conference on the Foundations of Software Engineering; June 23–28, 2025; Trondheim, Norwaybooktitle: 33rd ACM International Conference on the Foundations of Software Engineering (FSE Companion ’25), June 23–28, 2025, Trondheim, Norwaydoi: 10.1145/3696630.3731665isbn: 979-8-4007-1276-0/2025/06ccs: Human-centered computing Empirical studies in HCIccs: Software and its engineering Programming teamscopyright: none
1. Introduction
@@
Artificial Intelligence (AI) has revolutionized how we think about software engineering. According to the most recent StackOverflow Developer Survey, 76.7% of developers are already using or plan to incorporate AI-assisted software development assistants into their workflows (sta, [n. d.]). Developers use AI-assisted software tools in a variety of contexts and to complete a diversity of tasks, most often hoping for increases in their productivity (Coutinho et al., 2024).
@@
A significant contributor to, or deterrent from, developer productivity is the time they spend seeking information to support the completion of their tasks (Gonçalves et al., 2011; Meyer et al., 2019). Over the years, we have acquired an in-depth understanding of the information needs developers have (Ko et al., 2007; Liu et al., 2021; Phillips et al., 2012), the ways in which they attempt to meet their information needs (Fritz and Murphy, 2010), and the challenges they encounter in the process (LaToza and Myers, 2010). We also know that developers’ ability to meet information needs contributes to their ability to acquire and build expertise (Begel and Simon, 2008; Ford et al., 2017).
@@
Given the changing landscape in developer tooling, many studies have investigated the use and impact of AI-assisted tools on developer engagement and productivity (Ziegler et al., 2022; Cheng et al., 2024; Chinthapatla, 2024; Storey and Zagalsky, 2016). However, most of these studies focus on AI-assisted tool use in the context of understanding, generating, or modifying source code, providing little insight into the role AI plays in developer information seeking and its impact. To address this gap, we conducted a mixed methods study to answer the following research questions:
RQ1:
@@
To answer our research questions, we first administered a survey ($n=128$) to better understand to what extent developers engage with AI tools in this context. Building on those insights, we conducted a series of 17 interviews to further contextualize participants’ experiences. Our findings reveal that while AI tools can offer increased efficiency when seeking information, there are caveats to realizing the benefits, such as avoiding over-reliance and building the necessary expertise to appropriately and effectively use AI tooling to meet information needs.
2. Related Work
@@
The goal of our research is better understand information seeking with the availability of AI-assisted tools. Most relevant to our efforts are investigations into the information needs and seeking behaviors of developers and human-centric concerns regarding the use of AI tools in software development.
2.1. Information Needs of Developers
@@
Previous studies have examined the information needs of developers during the software development process. One of the first efforts centered on developer information needs was conducted by Ko et al. (Ko et al., 2007), where they observed developers in their work to better understand information needs of collocated software teams. Also aiming to gain a broad understanding of developer information needs is the work done by Fritz et al. (Fritz and Murphy, 2010), who conducted a series of 11 interviews to explore the use of information fragments to answer developers’ questions. On the flip side, LaToza et al. (LaToza and Myers, 2010) surveyed 179 professional developers regarding hard-to-answer questions about code to better understand information needs developers encounter challenges trying to fulfill.
@@
Some prior efforts are more focused, with an interest in understanding specific information needs. Breu et al. (Breu et al., 2010) analyzed information needs in bug reports by examining questions asked in six hundred bug reports from the Mozilla and Eclipse projects. Liu et al. (Liu et al., 2021) investigated API-related developer information needs on Stack Overflow by annotating Stack Overflow questions to APIs. Phillips et al. (Phillips et al., 2012) explored information needs for integration decisions in the release process of large-scale parallel development by interviewing seven release managers. With the emergence of AI-assisted tools, our builds on prior efforts by providing insights into the ways in which these tools have changed software practitioners’ information-seeking behaviors.
2.2. AI Tool Use in Software Engineering
Russo (Russo, 2024) identified factors that influence adoption of Generative AI tools in software engineering and found that compatibility with existing development workflows was the primary driver of AI tool adoption, rather than perceived usefulness or social factors. Others have studied the context and impact of integrating AI tools into software engineering processes, some of which we discuss below. Barke et al. (Barke et al., 2023) investigated how developers interact with AI programming assistants like GitHub Copilot. They classified interaction modes into two types: acceleration mode, where developers know the next steps and use Copilot to speed up their work, and exploration mode, where developers are unsure and use Copilot to discover possible solutions. Bird et al. (Bird et al., 2022) explored early user experiences with Copilot in pair-programming contexts, revealing that while the tool can provide useful initial code suggestions, it often requires additional review, hence shifting some of the workload from coding to code validation. Liang et al. (Liang et al., 2024) conducted a large-scale survey to investigate the usability of AI programming assistants, focusing on the successes and challenges developers face when using these tools.
Johnson et al. (Johnson et al., 2023) developed the PICSE framework, which identifies key factors influencing engineers’ trust and usage of traditional and AI-assisted software tools. Their framework was derived from interviews with software practitioners.
Given the concern regarding the impact of AI-assisted tool use on productivity, prior work has also investigated how we can measure developer productivity when using code completion tools like GitHub Copilot1(Ziegler et al., 2022).
Our work builds on these prior efforts by providing insights into how these developer interactions with AI tools influence not just task completion, but also their information-seeking behaviors and its impact on their productivity and skill development.
3. Methodology
@@
3.1. The Survey
We designed a 20-minute survey administered through Qualtrics to engage developers in our research. [^1] This research is approved under IRBNet #: 2163056-1. Our survey was divided into four sections. The first section asked a series of demographic and background questions, such as current job role and years of programming experience. The design of this section was heavily influenced by the annual Stack Overflow Developer Survey (sta, [n. d.]). The next section asked questions about the specific AI-assisted tools they use, the frequency and purpose of use, and rationale behind their usage. We also asked questions about their information seeking behaviors with and without AI assistance, including the advantages and disadvantages for each. To assess the impact of AI tools on development tasks, we also asked respondents to evaluate how these tools influenced their approach to tasks like debugging, testing, implementing, and planning. In the final section of the survey, we included questions regarding the role of AI tool in learning and integrating new technologies, such as the impact on learning curves and the potential pitfalls. We also gave respondents the opportunity at the end of the survey to express interest in a follow-up interview (Section 3.2). Prior to administering our survey, we piloted with 3 respondents. This helped ensure clarity and identify any issues before deployment. We excluded this data from our final dataset. The survey instrument is available in our supplemental materials (Anonymous, 2025).
@@ 3.1. The Survey
Given that we distributed our survey through social platforms, there is a heightened risk of receiving invalid responses (Griffin et al., 2021). To mitigate the potential for analyzing invalid data, we filtered out responses that were completed in under three minutes or incomplete. This initial filtering left us with 168 responses. We further excluded responses that contained irrelevant content in their open-ended answers, as this suggests they may not have been giving due consideration to the survey. This resulted in our final dataset, containing 128 valid responses, which we used to report our findings.
Table 1. Age Groups of Survey Respondents
@@
Table 2. Professional Programming Experience
@@
Table 3. Job Titles of Survey Respondents
@@
Table 4. AI-Assisted Tools Used by Survey Respondents
@@
In our final dataset of responses (128), the age of our survey respondents ranged from 18 to over 65, median being 25-34 [Table 1] with of 4 years of experience (Table 2). Prior research suggests that developers with less experience spend more time seeking information and learning as they transition from novices to experts (Begel and Simon, 2008) (Ford et al., 2017) (Li et al., 2020), making this demographic particularly valuable for our study. Majority of our survey respondents actively use AI tools (83.6%) such as ChatGPT and GitHub Copilot (Table 4) in their work, while 16.4% occasionally use them. In addition to formal education, a majority of our survey respondents also used books and online resources such as forums and courses to learn programming. We interviewed a subset of these respondents, depending on their interest and availability and report on their demographics in Table  5.
@@
To analyze our survey, we first mapped each survey question to our research questions. We categorized experience levels of participants into five different groups, as outlined in Table 2. For closed ended questions, we ran descriptive statistics and applied Fisher’s exact test with Bonferroni correction to identify any correlations. We also report on frequencies to supplement our findings. For open-ended responses, we used thematic summaries to report findings due to the small sample size of valid and useful responses. We only report significant findings in Section 4 .
3.2. The Interviews
To supplement our survey findings, we conducted interviews over Zoom to gather detailed insights into participants’ information-seeking practices, experiences with AI-assisted tools, and their impact on productivity and skill development. This helped us obtain a variety of perspectives and allowed for immediate follow-up questions for clarification.
Each interview began with a discussion of participants’ background in software development, their experiences with AI-assisted tools, and how these tools fit into their workflow. Topics covered included the impact of AI tools on productivity, skill development, and problem-solving strategies. We also asked questions regarding their roles and responsibilities, the typical activities in their jobs, and their reliance on other developers or software tools. This includes what AI tools they use, or avoid use of, when information seeking and their satisfaction with the interactions. We elicited specific examples of their use to better understand aspects such as workflow integration and problem solving approaches. Lastly, we inquired about the impact team dynamics and productivity.
We piloted the interview protocol following best practices with one author and two participants to identify and correct any issues. The final interview protocol is available in our supplemental materials (Anonymous, 2025). We continued interviewing participants until we reached theoretical saturation, where no new themes or insights emerged from the interview (Guest et al., 2006). Table 5 details the background of our interview participants.
Table 5. Interview Participant Background
@@ 3.2. The Interviews
To analyze our interview data, we used qualitative coding and thematic analysis, beginning with the creation of an initial codebook based on questions asked during the participant interviews. In the first iteration, all four authors independently labeled an interview transcript using the initial codebook and then collectively discussed their findings, which led to the emergence of a new set of codes. After agreement on the codes and how they should be used, we incorporated the emergent codes to finalize the codebook. Following this, we applied the final codebook (Anonymous, 2025) to label the remaining transcripts.
To ensure the rigor and validity of our efforts, we invited an external auditor to review a subset of raw coded segments from the interview transcripts along with the codebook. We asked the auditor to confirm that the inferences we made from the data were sound. We discussed and clarified any concerns and updated code descriptions when deemed necessary. After validating the codes, the first author reviewed all coded segments under each code and documented the overarching themes as they emerged. Finally, we mapped the codes to our research questions and iteratively identified any overlapping themes.
4. Results
The goal of our efforts are to uncover when, why, and how developers are using AI tools to support their information seeking, and their impact on developer productivity and skill development. Below we discuss our findings regarding the considerations developers may make when deciding to engage with AI tools to meet their information needs.
4.1. Information Seeking with AI tools (RQ1)
@@
Our findings suggest that developers are often using AI tools to support their information seeking Figure 1), with the majority reporting using AI tools at least half of the time in their work. However, according to the experiences reported in our interviews, their goals when doing so are most often broad and aimed at understanding the necessary considerations for completing their task.
@@
Figure 1. Frequency of AI Use for Information Seeking
Most participants reported relying on AI tools when trying to understand best practices, discover new libraries or solutions or explore trade-offs between different libraries and implementations.
Participants in our study also cited AI tools as being useful for identifying keywords for further online search and validation.
Another common use case among participants when information seeking is to use AI tools to recall previous knowledge or explain code. Participants mentioned value in using AI tools for synthesizing relevant information from documentation, which includes providing boilerplate code, pinpointing highly specific issues or information. One participant shared a concrete example of this
@@
UX Issues in AI-assisted Information Seeking
While participants in our study outlined numerous benefits to AI-assisted information seeking, concerns surfaced regarding the usability of AI tools. Most prominent were discussions regarding the non-prescriptive language used by AI tools. Rather than identifying critical requirements as mandatory, AI tools often present information using deferential suggestions that can undermine technical imperatives, as one participant explained:
“For something like code review, it’s a bit riskier because error handling is not just a nice-to-have; it’s a must-have. While the AI correctly identifies many issues, it frames them as improvements rather than necessities.”[P15]
Participants also highlighted how AI’s adaptive responses create unique challenges for technical information validation. Unlike static documentation or community forums where information remains consistent, participants observed that AI tools can shift their technical guidance based on user interactions: “It provided the correct information initially but later changed its stance to align with what I was saying. So now I don’t know when to trust it.” This inconsistency particularly undermines AI tools’ utility as authoritative technical references. Additionally, participants reported issues with inappropriate information density, noting that “Sometimes they tend to give out too much information, sometimes too little,” making it difficult to efficiently extract needed knowledge. We discuss more about these impacts on learning new technologies in Section 4.3.
@@
Figure 2. Traditional vs. AI preference for tasks
Task Legend: D* = Debugging, CA* = Code Analysis, R* = Refactoring, CC* = Code Comprehension, B* = Brainstorming, LNT* = Learning New Technologies.
AI-Generated Information Validation Process
Our findings reveal a contrast between traditional and AI-based information validation. Traditional sources rely on established mechanisms such as community vetting—creating standardized validation pathways. AI-generated content, however, lacks such mechanisms. Participants described not only the burden of validation but also the challenge of incomplete information. One participant’s experience illustrates this problem: “ChatGPT provided a partial answer that it scraped from Python docs regarding ArgParse library. I had to go to the actual docs to find it had a lot of missing context and prior steps.” This fragmented information delivery requires users to fill substantial knowledge gaps. Many employ cross-referencing between multiple AI systems to identify conflicts and inconsistencies, as one explained: it’s especially helpful to use multiple AI assistants” to find a whole pile of issues […] Hallucinations don’t seem to compound.”. The efficiency gains promised by AI-assisted information seeking are partially offset by these added validation demands, a challenge that current AI interfaces do little to address.
@@
The most common resource mentioned for information seeking without AI assistance was Stack Overflow, which participants cited using for context-specific tasks, such as resolving environment-specific issues, in the hopes that other developers have encountered similar scenarios. For GUI-related tasks, participants often preferred official documentation, with one participant noting how for Angular components, they would go there and search for the template” instead of using AI tools, as the official documentation provides proper solution[s] that [are] 100% correct.” Another source participants mentioned using was Reddit, where they would seek general guidance, such as tool recommendations, from specific subreddits. We also found that when it comes to tasks that require specific information, knowledge, or context, or when working with specialized or niche libraries, our participants sought support directly from other developers or support teams on contract.
@@
Regardless of the benefits, and perhaps due in part to the concerns, that come with using AI tools for seeking information, participants in our study highlighted several external factors that influence the impact and their use of AI tools in software development.
Participants noted that their usage of AI tools for information seeking is heavily influenced by the adoption and promotion of AI within the organization. Some organizations are proactive in adopting and promoting AI use, aiming to utilize these tools responsibly and in a controlled manner, while others feel “it’s not worth your time”. However, this proactive stance can sometimes lead to an overshadowing of traditional engineering practices, resulting in less investment in non-AI related professional development.